OpenStudy (anonymous):

Find the midpoint of the segment with the given endpoints. A(3, –5) and C(2, 9) (1 point) (–1, 5.5) (0.5, –7) (2.5, 2) (8, –7) @Feel-Like-A-Sir

OpenStudy (anonymous):

The answer is (2.5, 2).

OpenStudy (anonymous):

okay here's another one. Find the midpoint of AB- https://www.connexus.com/content/media/102877-3182010-93934-PM-1187492183/res00000/ppg/examview/images/mc012-2.jpg (3, 1) (1, 1) (10, 4) (4, 4)

OpenStudy (anonymous):

(1, 1)

OpenStudy (anonymous):

Doing these is quite simple. For the first one, I used the formula: (x1+x2)/2, (y1+y2)/2 So you put in the (3,-5) and the (2,9) in there. (3+2)/2, (-5+9)/2 5/2, 2 Then you divide 5/2 and get 2.5. Hence, you have the answer 2.5, 2.

OpenStudy (anonymous):

This is a big grade :) I'm glad I have your help.

OpenStudy (anonymous):

:) No problem.

OpenStudy (anonymous):

I need help with the one up there ^

OpenStudy (anonymous):

Yeah, I'm working on it. :3

OpenStudy (anonymous):

okay thank you :*

OpenStudy (anonymous):

Sorry, but I can't seem to figure this one out.. :s

OpenStudy (anonymous):

It's okay , we could come back to it later .. here's another one Find the length of the hypotenuse. Round to the nearest tenth if necessary. https://www.connexus.com/content/media/102877-3182010-93934-PM-1187492183/res00000/ppg/examview/images/mc014-1.jpg 6.9 ft 2.8 ft 5.7 ft 8 ft

OpenStudy (anonymous):

?

OpenStudy (anonymous):

One sec, I'm working on it. :P

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

Ingrid is making a quilt using squares that measure 5 in. on a side. What is the length of a diagonal of one of the quilt squares? Round to the nearest tenth. (1 point) 8.7 in. 7.1 in. 3.5 in. 14.2 in.

OpenStudy (anonymous):

You just use the formula a^2+b^2=c^2. 4^2 + 4^2 = c^2 16 + 16 = c^2 32 = c^2 To get rid of the "^2" on c, you have to square both sides. So, when you square 32, you get 5.66 which rounds up to 5.7. c = 5.7.

OpenStudy (anonymous):

^ That reply was for the hypotenuse question, btw.

OpenStudy (anonymous):

The answer is 7.1 in.

OpenStudy (anonymous):

Find the missing lengths in the triangle. Round to the nearest tenth if necessary. https://www.connexus.com/content/media/102877-3182010-93934-PM-1187492183/res00000/ppg/examview/images/mc016-1.jpg a = 7.5 cm, b = 10.6 cm a = 13 cm, b = 7.5 cm a = 7.5 cm, b = 13 cm a = 5 cm, b = 13 cm

OpenStudy (anonymous):

@Feel-Like-A-Sir

OpenStudy (anonymous):

Helpp ? :((